OpenStudy (anonymous):

Find the y-coordinate of the vertex for the following function: g(x) = x2 - 3x + 2

OpenStudy (anonymous):

we found the x coordinate right? it is \[\frac{3}{2}\] find the y coordinate by replacing x by \[\frac{3}{2}\] to see what you get

OpenStudy (anonymous):

you need the steps? it amounts to computing \[(\frac{3}{2})^2-3\times \frac{3}{2}+2\]

OpenStudy (anonymous):

we didn't find the x-coordinate. that was the axis of symmetry

OpenStudy (anonymous):

if you do it carefully you should get \[-\frac{1}{4}\]

OpenStudy (anonymous):

that is also the first coordinate of the vertex.

OpenStudy (anonymous):

oohh. ok thanks.

OpenStudy (anonymous):

how would i find the y-intercept?

OpenStudy (anonymous):

a bit confusing because when you write the axis of symmetry is \[x=\frac{3}{2}\] you are thinking of the vertical line given by the equation, but when you write the vertex you are thinking that \[x=\frac{3}{2}\] means x is the number \[\frac{3}{2}\] and so the vertex is \[(\frac{3}{2},-\frac{1}{4})\]

OpenStudy (anonymous):

find the y intercept by setting x = 0 to see what you get that is an easy one because if x = 0 then y = 2 pretty much by your eyeballs

OpenStudy (anonymous):

thanks a lot. this was a lot of help.